OpenStudy (1dfanatic):

A line passes through the point (4,-1) and has a slope of 9/2. Find the equation of the line

OpenStudy (isaidavila):

y2-y1=m(x2-x1) y-(-1)=9/2(x-4) y+1=9/2(x-4)

OpenStudy (isaidavila):

Hope it helps

OpenStudy (anonymous):

I thought it was y-y1=m(x-x1)

OpenStudy (anonymous):

And the y2-y1 and x2-x1 was slope

OpenStudy (isaidavila):

It's the same.

OpenStudy (anonymous):

Oh okay :)

OpenStudy (1dfanatic):

Im kinda confused

OpenStudy (isaidavila):

Don't worry about the number of y and x

OpenStudy (anonymous):